# A Note on Estimation Error Bound and Grouping Effect of Transfer Elastic Net

### v1

@@ 1 Introduction
Transfer learning refers to a set of techniques to utilize the knowledge acquired from solving source problems to solve another target problem. Recently, Takada and Fujisawa, (2020) proposed the Transfer Lasso, which uses the $\ell_{1}$ norm to transfer the information of the source estimates to the target problem, which encourages the sparsity in the estimates and the sparsity in the changes of the estimates. Tomo and Nakaki, (2024) proposed the Transfer Elastic Net to address high-dimensional features in the field of bioinformatics, which employs a combination of $\ell_{1}$ and $\ell_{2}$ norms to facilitate knowledge transfer. Although Colangelo, (2021) proposed a similar estimator, they do not use their $\ell_{2}$ norm for knowledge transfer. In the following section, we define the estimator of the Transfer Elastic Net proposed in Tomo and Nakaki, (2024).
@@
where $\lambda\in[0,+\infty)$, $\alpha\in[0,1]$, and $\rho\in[0,1]$ are the tuning parameters, which control the intensity of the regularization, the balance of the $\ell_{1}$ and $\ell_{2}$ norm, and the extent of transferring information, respectively. The estimator $\hat{\boldsymbol{\beta}}_{\mathrm{TENet}}$ is defined as the minimizer of the loss function:
@@ 2 Estimation Error Bound
Let $\boldsymbol{\Delta}=\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}$ and $\boldsymbol{\Delta}_{\alpha}=\boldsymbol{\Delta}-\alpha\tilde{\boldsymbol{\beta}}$ ($={(1-\alpha)\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta}}^{*}$). Let $\boldsymbol{v}_{S}$ denote the subvector of $\boldsymbol{v}\in\mathbb{R}^{p}$ obtained by extracting the entries indexed by the index set $S$. We assume that $\boldsymbol{\beta}_{S}^{*}\in\mathbb{R}^{s}$ is non-zero in all its elements, and $\boldsymbol{\beta}_{S^{c}}^{*}\in\mathbb{R}^{p-s}$ is the zero vector. We establish the following assumptions to derive an estimation error bound for the Transfer Elastic Net.
@@
The error terms $\varepsilon_{1},\ldots,\varepsilon_{n}$ satisfies the conditions below.
@@
From Theorem 1 and Corollary 2, we establish the propositions that provide the comparison of the bounds. We assume that the source problem is highly related to the target problem and that $\tilde{\boldsymbol{\beta}}=\boldsymbol{\beta}^{*}$. Then, we prove the following proposition providing a relationship between the error bounds of the Transfer Elastic Net and the ordinary Elastic Net.
@@
We then consider the Transfer Lasso. The following proposition shows a sufficient condition under which $U_{\mathrm{TENet}}$ is guaranteed to be smaller than $U_{\mathrm{TLasso}}$.
@@ 3 Grouping Effect
The grouping effect of the Elastic Net ensures that the estimates of the coefficients corresponding to highly correlated predictors have a small difference (Zou and Hastie,, 2005; Zhou,, 2013). We show that the Transfer Elastic Net retains a similar property. Let $r_{jk}$ denote the correlation coefficient between the $j$th and $k$th column of $\mathbf{X}$. Then we prove the following theorem.
@@
This theorem suggests that the estimates corresponding to the strongly correlated variables exhibit a small difference if the corresponding source estimates have a small difference.
Now suppose that the source estimate $\tilde{\boldsymbol{\beta}}$ was obtained by applying the ordinary Elastic Net to the data with the response ${\boldsymbol{y}^{\prime}}\in\mathbb{R}^{m}$ and the predictors ${\mathbf{X}}^{\prime}\in\mathbb{R}^{m\times p}$, with $\sum_{i=1}^{n}y^{\prime}_{i}=0$, $\sum_{i=1}^{n}X_{i,j}^{\prime}=0$, and $\sum_{i=1}^{n}{X_{i,j}^{\prime}}^{2}/n=1$ for $j=0,\ldots,p-1$. Therefore, we assume that $\tilde{\boldsymbol{\beta}}$ is defined as:
@@
$\left|\tilde{{\beta}}_{j}-\tilde{{\beta}}_{k}\right|$ $\leq\frac{\|\tilde{\boldsymbol{y}}\|_{2}\sqrt{2n(1-\tilde{r}_{jk})}}{2n\lambda^{\prime}(1-\rho^{\prime})},$
@@
Acknowledgement
The authors thank Takayuki Kawashima for valuable comments.
@@ Appendix C Proof of Proposition 4
Appendix D Proof of Theorem 5

### v2

@@ 1 Introduction
Transfer learning refers to a set of techniques to utilize the knowledge acquired from solving source problems to solve another target problem. Recently, Takada and Fujisawa, (2020) proposed the Transfer Lasso, which uses the $\ell_{1}$ norm to transfer the information of the source estimates to the target problem, which encourages the sparsity in the estimates and the sparsity in the changes of the estimates. Tomo and Nakaki, (2024) proposed the Transfer Elastic Net to address high-dimensional features in the field of bioinformatics, employing a combination of $\ell_{1}$ and $\ell_{2}$ norms with the expectation of mitigating instability caused by correlations among variables. Although Colangelo, (2021) proposed a similar estimator, they do not use their $\ell_{2}$ norm for knowledge transfer. In the following section, we define the estimator of the Transfer Elastic Net proposed in Tomo and Nakaki, (2024).
@@
where $\lambda\in[0,+\infty)$, $\alpha\in[0,1]$, and $\rho\in[0,1]$ are the tuning parameters, which control the intensity of the regularization, the balance of the $\ell_{1}$ and $\ell_{2}$ norms, and the extent of transferring information, respectively. The estimator $\hat{\boldsymbol{\beta}}_{\mathrm{TENet}}$ is defined as the minimizer of the loss function:
@@ 2 Estimation Error Bound
Let $\boldsymbol{\Delta}=\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta}^{*}$ and $\boldsymbol{\Delta}_{\alpha}=\boldsymbol{\Delta}-\alpha\tilde{\boldsymbol{\beta}}$ ($={(1-\alpha)\tilde{\boldsymbol{\beta}}-\boldsymbol{\beta}}^{*}$). Let $\boldsymbol{v}_{S}$ denote the subvector of $\boldsymbol{v}\in\mathbb{R}^{p}$ obtained by extracting the entries indexed by the index set $S$. We assume that $\boldsymbol{\beta}_{S}^{*}\in\mathbb{R}^{s}$ is non-zero in all its elements, and $\boldsymbol{\beta}_{S^{c}}^{*}\in\mathbb{R}^{p-s}$ is the zero vector. We assume the following conditions to derive an estimation error bound for the Transfer Elastic Net.
@@
The error terms $\varepsilon_{1},\ldots,\varepsilon_{n}$ satisfy the conditions below.
@@
From Theorem 1 and Corollary 2, we establish the propositions that provide the comparison of the bounds. We assume that the source problem is highly related to the target problem and that $\tilde{\boldsymbol{\beta}}=\boldsymbol{\beta}^{*}$. Then, we prove the following proposition providing a relationship between the error bounds of the Transfer Elastic Net and the ordinary Elastic Net for any $\alpha$ and the same $\lambda$ and $\rho$.
@@
We then consider the Transfer Lasso. The following proposition shows a sufficient condition under which $U_{\mathrm{TENet}}$ is guaranteed to be smaller than $U_{\mathrm{TLasso}}$ for any $\rho$ and the same $\lambda$ and $\alpha$.
@@
Now we discuss the conditions under which the generalized restricted eigenvalue condition is satisfied. From the result of Raskutti et al., (2010), the ordinary restricted eigenvalue condition for the Lasso is satisfied with high probability when predictors are independent and identically distributed (i.i.d.) samples from a Gaussian distribution with a covariance matrix that satisfies the condition. For the generalized restricted eigenvalue condition, we prove the following proposition.
Proposition 5.
Suppose $02\alpha\rho-c-\rho>$ 0. Suppose $\boldsymbol{X}_{1},\ldots,\boldsymbol{X}_{n}$ are $i.i.d.$ samples from $\mathcal{N}(\boldsymbol{0},\mathbf{\Sigma})$ with $\mathbf{\Sigma}\in\mathbb{R}^{p\times p}$ and the following inequality holds for some $\gamma>0$:
${\boldsymbol{v}^{\top}\mathbf{\Sigma}\boldsymbol{v}}>\gamma\|\boldsymbol{v}\|_{2}^{2},\quad\text{for all}\,\boldsymbol{v}\in\mathcal{B}(\alpha,\rho,c,\boldsymbol{\Delta}).$
Under these conditions, there exist universal constants $c^{\prime},c^{\prime\prime},c^{\prime\prime\prime}>0$ such that if the sample size satisfies
$n$ $>\frac{c^{\prime\prime\prime}M}{\gamma}{\left(\frac{2\alpha\rho\sqrt{s}+2(1-\rho)\left\|\boldsymbol{\Delta}_{\alpha}\right\|_{2}}{2\alpha\rho-c-\rho}\right)^{2}}\log p,$
where $M:=\max_{j=1,\ldots,p}\Sigma_{jj}$, then, we have
$\frac{1}{n}{\boldsymbol{v}^{\top}\mathbf{X}^{\top}\mathbf{X}\boldsymbol{v}}>\frac{\gamma}{64}{\|\boldsymbol{v}\|_{2}^{2}}\quad\text{for all}\,\boldsymbol{v}\in\mathcal{B}(\alpha,\rho,c,\boldsymbol{\Delta}),$
with probability at least $1-c^{\prime\prime}\exp(-c^{\prime}n)$.
This proposition suggests that a broad class of Gaussian matrices satisfies the generalized restricted eigenvalue condition for $\mathcal{B}=\mathcal{B}(\alpha,\rho,c,\boldsymbol{\Delta})$ with high probability.
@@ 3 Grouping Effect
The grouping effect of the Elastic Net ensures that the estimates of the coefficients corresponding to highly correlated predictors have a small difference (Zou and Hastie,, 2005; Zhou,, 2013). We show that the Transfer Elastic Net retains a similar property. Let $r_{jk}$ denote the correlation coefficient between the $j$th and $k$th column of $\mathbf{X}$. We write $\hat{\boldsymbol{\beta}}=\hat{\boldsymbol{\beta}}_{\mathrm{TENet}}$. Then we prove the following theorem.
@@
This theorem suggests that the estimates corresponding to strongly correlated variables exhibit a small difference if the corresponding source estimates have a small difference or $\alpha$ is close to $1$. Although the conditions of the theorem involve estimators, a similar condition is presented in the results of Zou and Hastie, (2005).
Now suppose that the source estimate $\tilde{\boldsymbol{\beta}}$ was obtained by applying the ordinary Elastic Net to the data with the response ${\boldsymbol{y}^{\prime}}\in\mathbb{R}^{m}$ and the predictors ${\mathbf{X}}^{\prime}\in\mathbb{R}^{m\times p}$, with $\sum_{i=1}^{m}y^{\prime}_{i}=0$, $\sum_{i=1}^{m}X_{i,j}^{\prime}=0$, and $\sum_{i=1}^{m}{X_{i,j}^{\prime}}^{2}/m=1$ for $j=0,\ldots,p-1$. Therefore, we assume that $\tilde{\boldsymbol{\beta}}$ is defined as:
@@
$\left|\tilde{{\beta}}_{j}-\tilde{{\beta}}_{k}\right|$ $\leq\frac{\|\tilde{\boldsymbol{y}}\|_{2}\sqrt{2m(1-\tilde{r}_{jk})}}{2m\lambda^{\prime}(1-\rho^{\prime})},$
@@
Remark 2.
When the regularization term is defined as
$R(\boldsymbol{\beta},\tilde{\boldsymbol{\beta}};\alpha,\rho_{1},\rho_{2})$ $:=\alpha\left\{\rho_{1}\|\boldsymbol{\beta}\|_{1}+(1-\rho_{1})\|\boldsymbol{\beta}\|_{2}^{2}\right\}+(1-\alpha)\left\{\rho_{1}\|\boldsymbol{\beta}-\tilde{\boldsymbol{\beta}}\|_{1}+\rho_{2}\|\boldsymbol{\beta}-\tilde{\boldsymbol{\beta}}\|_{2}^{2}\right\},$
and we set $\rho_{2}=0$, then we have
$\left|\hat{{\beta}}_{j}-\hat{{\beta}}_{k}\right|$ $\leq Z\sqrt{1-r_{jk}},\quad Z=\sqrt{\frac{\left\|\boldsymbol{y}\right\|_{2}^{2}+2n\lambda(1-\alpha)\rho_{1}\|\tilde{\boldsymbol{\beta}}\|_{1}}{2n\alpha^{2}\lambda^{2}(1-\rho_{1})^{2}}},$
if $\hat{{\beta}}_{j}\hat{{\beta}}_{k}>0$ and $(\hat{{\beta}}_{j}-\tilde{{\beta}}_{j})(\hat{{\beta}}_{k}-\tilde{{\beta}}_{k})>0$, according to the same derivation. This result suggests that the estimates corresponding to strongly correlated variables exhibit a small difference, regardless of $\alpha\neq 0$, $\rho_{1}\neq 1$, and the difference of $\tilde{\beta}_{j}$ and $\tilde{\beta}_{k}$ if the knowledge transfer is achieved solely through the $\ell_{1}$ norm term.
Acknowledgements
The authors thank Takayuki Kawashima for providing valuable comments on an early draft of this manuscript. The authors are also deeply grateful to Masaaki Takada for helpful discussions and suggestions regarding aspects that were insufficiently addressed.
@@ Appendix C Proof of Proposition 4
Appendix D Proof of Proposition 5
We first introduce the following theorem from Raskutti et al., (2010):
Theorem 9 (Theorem 1 of Raskutti et al., (2010)).
For any Gaussian random design $\mathbf{X}\in\mathbb{R}^{n\times p}$ with i.i.d. $\mathcal{N}(0,\mathbf{\Sigma})$ rows, there are universal positive constants $c^{\prime},\,c^{\prime\prime}$ such that
$\frac{\|\mathbf{X}\boldsymbol{v}\|_{2}}{\sqrt{n}}$ $\geq\frac{1}{4}\left\|\mathbf{\Sigma}^{1/2}\boldsymbol{v}\right\|_{2}-9M\sqrt{\frac{\log p}{n}}\|\boldsymbol{v}\|_{1}\quad\text{for all }\,\boldsymbol{v}\in\mathbb{R}^{p},$
where $M:=\max_{j=1,\ldots,p}\Sigma_{jj}$, with probability at least $1-c^{\prime\prime}\exp(-c^{\prime}n)$.
Then, using this theorem, we prove Proposition 5 as follows.
Proof of Proposition 5.
From the definition, all $\boldsymbol{v}\in\mathcal{B}(\alpha,\rho,c,\boldsymbol{\Delta})$ satisfies
$(\alpha\rho-c)\|\boldsymbol{v}_{S^{c}}\|_{1}+(1-\alpha)\rho\|\boldsymbol{v}-\boldsymbol{\Delta}\|_{1}$ $\leq(\alpha\rho+c)\|\boldsymbol{v}_{S}\|_{1}+(1-\alpha)\rho\|\boldsymbol{\Delta}\|_{1}+2(1-\rho)\|\boldsymbol{\Delta}_{\alpha}\|_{2}\|\boldsymbol{v}\|_{2}.$ (3)
From the triangle inequality $\|\boldsymbol{v}-\boldsymbol{\Delta}\|_{1}\geq\|\boldsymbol{\Delta}\|_{1}-\|\boldsymbol{v}\|_{1}$, it holds that
$(\alpha\rho-c)\|\boldsymbol{v}_{S^{c}}\|_{1}+(1-\alpha)\rho\left(\|\boldsymbol{\Delta}\|_{1}-\|\boldsymbol{v}\|_{1}\right)$ $\leq(\alpha\rho+c)\|\boldsymbol{v}_{S}\|_{1}+(1-\alpha)\rho\|\boldsymbol{\Delta}\|_{1}+2(1-\rho)\|\boldsymbol{\Delta}_{\alpha}\|_{2}\|\boldsymbol{v}\|_{2}.$
Then, using $\|\boldsymbol{v}\|_{1}=\|\boldsymbol{v}_{S}\|_{1}+\|\boldsymbol{v}_{S^{c}}\|_{1}$, we have
$(2\alpha\rho-c-\rho)\|\boldsymbol{v}_{S^{c}}\|_{1}$ $\leq(\rho+c)\|\boldsymbol{v}_{S}\|_{1}+2(1-\rho)\|\boldsymbol{\Delta}_{\alpha}\|_{2}\|\boldsymbol{v}\|_{2}.$
Consequently, it holds that
$(2\alpha\rho-c-\rho)\|\boldsymbol{v}\|_{1}$ $\leq 2\alpha\rho\|\boldsymbol{v}_{S}\|_{1}+2(1-\rho)\|\boldsymbol{\Delta}_{\alpha}\|_{2}\|\boldsymbol{v}\|_{2}.$
By the norm inequality $\|\boldsymbol{v}_{S}\|_{1}\leq\sqrt{s}\|\boldsymbol{v}_{S}\|_{2}\leq\sqrt{s}\|\boldsymbol{v}\|_{2}$, we obtain
$(2\alpha\rho-c-\rho)\|\boldsymbol{v}\|_{1}$ $\leq\left(2\alpha\rho\sqrt{s}+2(1-\rho)\|\boldsymbol{\Delta}_{\alpha}\|_{2}\right)\|\boldsymbol{v}\|_{2}.$
Then, using $02\alpha\rho-c-\rho>$ 0, we have
$\|\boldsymbol{v}\|_{1}$ $\leq\frac{2\alpha\rho\sqrt{s}+2(1-\rho)\|\boldsymbol{\Delta}_{\alpha}\|_{2}}{2\alpha\rho-c-\rho}\|\boldsymbol{v}\|_{2}.$ (4)
From Theorem 9 and (4), for all $\boldsymbol{v}\in\mathcal{B}(\alpha,\rho,c,\boldsymbol{\Delta})$, it holds that
$\frac{\|\mathbf{X}\boldsymbol{v}\|_{2}}{\sqrt{n}}$ $\geq\frac{1}{4}\left\|\mathbf{\Sigma}^{1/2}\boldsymbol{v}\right\|_{2}-9M\sqrt{\frac{\log p}{n}}\cdot\frac{2\alpha\rho\sqrt{s}+2(1-\rho)\|\boldsymbol{\Delta}_{\alpha}\|_{2}}{2\alpha\rho-c-\rho}\|\boldsymbol{v}\|_{2},$
with probability at least $1-c^{\prime\prime}\exp(-c^{\prime}n)$. Using $\left\|\mathbf{\Sigma}^{1/2}\boldsymbol{v}\right\|_{2}\geq\sqrt{\gamma}\|\boldsymbol{v}\|_{2}$, we obtain
$\frac{\|\mathbf{X}\boldsymbol{v}\|_{2}}{\sqrt{n}}$ $\geq\left(\frac{\sqrt{\gamma}}{4}-9M\sqrt{\frac{\log p}{n}}\cdot\frac{2\alpha\rho\sqrt{s}+2(1-\rho)\|\boldsymbol{\Delta}_{\alpha}\|_{2}}{2\alpha\rho-c-\rho}\right)\|\boldsymbol{v}\|_{2}.$
Therefore, if the sample size $n$ satisfies
$n$ $>\frac{(72M)^{2}}{\gamma}\left(\frac{2\alpha\rho\sqrt{s}+2(1-\rho)\|\boldsymbol{\Delta}_{\alpha}\|_{2}}{2\alpha\rho-c-\rho}\right)\log p,$
then we have
$\frac{\|\mathbf{X}\boldsymbol{v}\|_{2}}{\sqrt{n}}$ $\geq\frac{\sqrt{\gamma}}{8}\|\boldsymbol{v}\|_{2},$
with probability at least $1-c^{\prime\prime}\exp(-c^{\prime}n)$. This completes the proof. ∎
Appendix E Proof of Theorem 6